OpenStudy (anonymous):

\3{^-x{^2}} = 7\ Use common log & two decimal places. Why does book say "no solution?"

OpenStudy (anonymous):

The result is 0.43429

OpenStudy (anonymous):

Is the book wrong, then? (that's fine... I'm just confused!)

OpenStudy (anonymous):

I guess it is

OpenStudy (anonymous):

go figure... Thanks very much! :-)

OpenStudy (mertsj):

\[3^{-x ^{2}}=7\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[-x ^{2}\ln 3=\ln 7\] \[-x^2=\frac{\ln 7}{\ln 3}=1.77\] \[x^2=-1.77\] Now do you see why there is no solution? x^2 cannot be negative

OpenStudy (anonymous):

Ok. That was the answer I came up with, too. Thank you much for this! :-)

OpenStudy (mertsj):

yw